# Multi-Agent Code Verification with Compound Vulnerability Detection

### v1

@@ Abstract
[…]
Keywords: Multi-agent systems, Code verification, LLM-generated code, Compound vulnerabilities
@@ 1 Introduction
Our Approach. We built CodeX-Verify, a system that runs four specialized agents in parallel: Correctness (logic errors, edge cases, exception handling), Security (OWASP Top 10, CWE patterns, secrets), Performance (algorithmic complexity, resource leaks), and Style (maintainability, documentation). Each agent looks for different bug types. We prove that combining agents finds more bugs than any single agent when the agents detect different problems: $I(A_{1},A_{2},A_{3},A_{4};B)>\max_{i}I(A_{i};B)$. Measuring how often our agents agree shows correlation $\rho=0.05$$0.25$, confirming they catch different bugs.
[…]
@@
    Mathematical proof that combining agents with different detection patterns finds more bugs than any single agent, measured by mutual information: $I(A_{1},A_{2},A_{3},A_{4};B)>\max_{i}I(A_{i};B)$. Agent correlation of $\rho=0.05$$0.25$ confirms they catch different bugs.
@@
    Formalization of compound vulnerability risk: when code has multiple security holes, the risk multiplies rather than adds. SQL injection (risk 10) plus hardcoded credentials (risk 10) creates risk 300, not 20, using amplification factors $\alpha\in\{1.5,2.0,2.5,3.0\}$ from attack graph theory.
@@
    Testing all 15 agent combinations on 99 samples shows multi-agent improves accuracy by 39.7 percentage points over single agents, with diminishing returns (+14.9pp, +13.5pp, +11.2pp) for each added agent.
@@
    Dataset of 99 code samples with verified labels covering 16 bug categories, achieving 76.1% TPR with 68.7% accuracy ($\pm$9.1% CI), released open-source.
@@ 2.1 LLM Code Generation and Verification
Wang and Zhu wang2024metamorphic propose Meta Prompt Testing: generate code variants with paraphrased prompts and detect bugs by checking if outputs differ. This achieves 75% TPR with 8.6% FPR on HumanEval. It requires test execution infrastructure and misses security vulnerabilities (SQL injection produces consistent outputs despite being exploitable) and quality issues that don’t affect outputs. AutoReview autoreview2025 uses three LLM agents (detector, locator, repairer) to find security bugs, improving F1 by 18.72% on ReposVul, but only checks security, not correctness or performance. We differ by: (1) proving mathematically why multi-agent works (information theory), (2) detecting compound vulnerabilities (multiple bugs amplifying risk), and (3) testing all 15 agent combinations to validate the architecture.
@@ 2.3 Static Analysis and Vulnerability Detection
We extend static analysis by: (1) coordinating multiple agents that check different bug types, (2) modeling how multiple vulnerabilities amplify risk, and (3) proving mathematically why combining analyzers works better.
@@ 2.4 Ensemble Learning and Information Theory
Sheyner et al. sheyner2002attack model multi-step network exploits as attack graphs (directed graphs of vulnerability chains). Later work bayesian2018attack adds Bayesian risk and CVSS scoring cvss2024. But attack graphs focus on network vulnerabilities (host compromise, privilege escalation), not code vulnerabilities. We adapt attack graph theory to code, modeling how SQL injection plus exposed credentials creates exponentially more risk ($\alpha>1$) than either alone.
@@ 3 Theoretical Framework
We prove why multi-agent code verification beats single-agent approaches, derive sample size requirements, and formalize compound vulnerability detection. Section 6 tests all theoretical predictions.
@@ 3.2 Why Multi-Agent Works
For agents with conditionally independent observations given code $c\sim\mathcal{C}$, the mutual information between combined agent observations and bug presence strictly exceeds that of any single agent:
$I(A_{1},A_{2},A_{3},A_{4};B)>\max_{i\in\{1,2,3,4\}}I(A_{i};B)$ (2)
whenever agents observe non-redundant bug patterns, i.e., $I(A_{i};B\mid A_{1},\ldots,A_{i-1})>0$ for some $i$.
@@
By the chain rule of mutual information:
$I(A_{1},A_{2},A_{3},A_{4};B)$ $=I(A_{1};B)+I(A_{2};B\mid A_{1})$
$\quad+I(A_{3};B\mid A_{1},A_{2})+I(A_{4};B\mid A_{1},A_{2},A_{3})$ (3)
Each term $I(A_{i};B\mid A_{1},\ldots,A_{i-1})\geq 0$ by definition, with equality only when $A_{i}$ adds no information beyond earlier agents (perfect redundancy).
Our agents check different bug types:
  • $A_{1}$ (Correctness): Logic errors, exception handling, edge case coverage
  • $A_{2}$ (Security): Injection vulnerabilities, hardcoded secrets, unsafe operations
  • $A_{3}$ (Performance): Algorithmic complexity, scalability, resource leaks
  • $A_{4}$ (Style): Maintainability metrics, documentation quality
These are different bug categories: a logic error ($A_{1}$) is distinct from SQL injection ($A_{2}$) and from $O(n^{3})$ complexity ($A_{3}$). So conditional information terms are positive:
$I(A_{1},A_{2},A_{3},A_{4};B)\geq I(A_{1};B)+\Delta_{2}+\Delta_{3}+\Delta_{4}$ (4)
where $\Delta_{i}=I(A_{i};B\mid A_{1},\ldots,A_{i-1})>0$ is agent $i$’s marginal contribution. ∎
Measured Results. Section 6.2 tests this by measuring agents alone and combined. Single agents get 17.2% to 75.9% accuracy, while 4 agents together get 72.4%, showing they complement each other. Measuring how often agents agree gives $\rho=0.05$$0.25$ (Table 3), confirming they detect different bugs. Progressive improvement (32.8% $\to$ 47.7% $\to$ 61.2% $\to$ 72.4% for 1, 2, 3, 4 agents) matches the prediction.
3.3 Ensemble Optimality and Diminishing Returns
Theorem 2 (Sublinear Information Gain).
When agents are ordered by decreasing individual performance, the marginal information gain from adding the $k$-th agent satisfies:
$\Delta I_{k}=I(A_{k};B\mid A_{1},\ldots,A_{k-1})\leq\Delta I_{k-1}$ (5)
in expectation, i.e., marginal contributions decrease monotonically.
Proof Sketch.
By data processing inequality, $I(A_{k};B\mid A_{1},\ldots,A_{k-1})$ decreases as we condition on more agents (more information already captured). When agents are ordered by performance, later agents add less. For any split into selected agents $\mathcal{A}$ and remaining agents $\mathcal{A}^{c}$, picking the best remaining agent yields monotonically decreasing gains. ∎
Corollary 1 (Optimal Agent Count). Optimal agent count $n^{*}$ is where marginal gain equals marginal cost. Our measurements suggest $n^{*}=4$: adding agents 2, 3, 4 yields +14.9pp, +13.5pp, +11.2pp (Section 6.2, Table 2). Extrapolating predicts agent 5 would add ¡10pp while increasing overhead, so 4 agents is near-optimal.
@@ 3.4 Weighted Aggregation
Proposition 3 (Optimal Weight Selection).
For agents with accuracies $p_{i}$ and correlations $\rho_{ij}$, approximately optimal weights are:
$w_{i}^{*}\propto p_{i}\cdot(1-\bar{\rho}_{i})\cdot\gamma_{i}$ (6)
where $\bar{\rho}_{i}=\frac{1}{n-1}\sum_{j\neq i}\rho_{ij}$ is agent $i$’s average correlation with others, and $\gamma_{i}$ is domain-specific criticality.
Higher-accuracy agents get higher weight ($p_{i}$), but weight decreases if the agent is redundant (high $\bar{\rho}_{i}$). The criticality term $\gamma_{i}$ captures asymmetric costs: security bugs block deployment more than style issues, justifying higher security weight despite lower accuracy.
@@
3.5 Compound Vulnerability Theory
An attack graph for code $c$ is $G_{c}=(V,E,\alpha)$ where:
  • $V\subseteq\mathcal{V}$ is the set of detected vulnerabilities
  • $E\subseteq V\times V$ represents exploitable chains: $(v_{i},v_{j})\in E$ if exploiting $v_{i}$ enables exploiting $v_{j}$
  • $\alpha:E\to\mathbb{R}^{+}$ quantifies how much vulnerabilities amplify each other
Theorem 4 (Exponential Risk Amplification).
For vulnerabilities $(v_{i},v_{j})\in E$ forming an attack chain, compound risk satisfies:
$\text{Risk}(v_{i}\cup v_{j})=\text{Risk}(v_{i})\times\text{Risk}(v_{j})\times\alpha(v_{i},v_{j})$ (7)
where $\alpha(v_{i},v_{j})>1$ represents the synergistic exploitation advantage unavailable to either vulnerability individually.
Proof.
Attack success with compound vulnerability:
[…]
P(chain succeeds)P($\text${chain succeeds}) captures the attacker’s ability to combine vulnerabilities. SQL injection alone gives limited access; credentials alone can’t be used. Together they enable full database access.
[…]
$\text{Risk}_{\text{compound}}=\text{Risk}(v_{i})\times\text{Risk}(v_{j})\times\alpha(v_{i},v_{j})$ (9)
Traditional models add risks: $R_{\text{linear}}=\sum_{i}\text{Risk}(v_{i})$. For SQL injection (risk 10) plus credentials (risk 10), this gives total risk 20. Our model with $\alpha=3.0$ gives:
$R_{\text{compound}}=10+10+(10\times 10\times 3.0-10-10)=300$ (10)
[…]
We set $\alpha=3.0$ for SQL injection + credentials, 2.0 for code execution + imports, 1.8 for complexity + inefficiency, calibrated from CVSS scores cvss2024.
@@ 3.7 Agent Selection
3.8 Ensemble Error Reduction
Proposition 7 (Ensemble Accuracy).
For $n$ classifiers with accuracy $p$ and correlation $\rho$:
$p_{\text{ensemble}}\approx p+\frac{(1-p)\cdot p\cdot(1-2p)}{1+(n-1)\rho}\cdot\sqrt{n}$ (16)
Improvement increases with $n$, decreases with $\rho$.
Uncorrelated errors ($\rho\to 0$): when $A_{1}$ misses a bug, $A_{2}$ catches it. Correlated ($\rho\to 1$): all miss the same bugs. Our $\rho=0.05$$0.25$ enables substantial gains.
@@ 3.10 Theory Summary
Compound $\alpha>1$ improves detection Compound detection active Yes
@@ 4.1 Architecture
CodeX-Verify runs a pipeline: code input $\to$ parallel agent execution $\to$ result aggregation $\to$ compound detection $\to$ deployment decision (Figure 1).
@@
              (Weighted combination,
           compound vulnerability detection)
@@
Design: (1) Agents check different bug types ($\rho=0.05$$0.25$ correlation). (2) Run in parallel via asyncio, ¡200ms latency. (3) Combine weighted scores, detect compounds, make decision.
@@ 4.2 Agent Specializations
Compound detection finds dangerous pairs:
1:procedure DetectCompounds($V$)
2:  $C\leftarrow\emptyset$
3:  for $(v_{i},v_{j})\in V\times V$ where $i<j$ do
4:   if $(v_{i}.\text{type},v_{j}.\text{type})\in E$ then
5:     $\alpha\leftarrow\textsc{GetAmplification}(v_{i},v_{j})$
6:     $\text{risk}\leftarrow\text{Risk}(v_{i})\times\text{Risk}(v_{j})\times\alpha$
7:     if $\text{risk}>\text{threshold}$ then
8:      $C\leftarrow C\cup\{(v_{i},v_{j},\text{risk})\}$
9:     end if
10:   end if
11:  end for
12:  return $C$
13:end procedure
Complexity: $O(|V|^{2})$, acceptable for $|V|<20$ per sample.
@@ 4.3 Aggregation
Agents run in parallel via asyncio (150ms max vs. 450ms sequential). Aggregation: collect issues, adjust severities by context, detect compounds (Algorithm 1), merge duplicates, compute $S_{\text{sys}}=\sum_{i}w_{i}S_{i}$, apply decision thresholds.
Compound pairs: (SQL injection, credentials), (code execution, dangerous import), (pickle.loads, network request), (complexity $O(n^{2})$, algorithm inefficiency). When both detected, amplify risk by $\alpha$ and flag CRITICAL.
@@ 4.4 Decision Logic
2:  if critical or compound vulnerabilities then
@@ 5.3 Ablation Design
We test all 15 combinations: single agents (4), pairs (6), triples (4), full system (1). Each tested on all 99 samples. Hypothesis: Theorem 1 predicts multi-agent beats single when correlation is low, with diminishing returns (Theorem 2). Marginal contribution: $\Delta_{i}=\mathbb{E}[\text{Acc}(\text{with }A_{i})]-\mathbb{E}[\text{Acc}(\text{without }A_{i})]$.
@@ 6.1 Main Evaluation Results
Different methodology (test-based vs. static) and dataset (HumanEval vs. ours).
$^{***}p<0.001$, $^{**}p<0.01$, $^{*}p<0.05$ (McNemar’s test, Bonferroni-corrected).
@@ 6.5 Real-World Validation on Claude Patches
Found 4 compounds: SQL + credentials (2, $\alpha=3.0$, risk 300), execution + import (1, $\alpha=2.0$, risk 200), complexity + inefficiency (1, $\alpha=1.8$, risk 180). All 4 flagged correctly (100%). Traditional additive: risk 20 (HIGH). Ours: risk 300 (CRITICAL, auto-blocks).
@@ 7.4 Limitations
Sample Size. $n=99$ gives ±9.1% confidence intervals, wider than ideal. PAC bound (Theorem 5) suggests $n\geq 127$ for $\epsilon=0.15$, so we’re below optimal. But our perfect labels (100% verified) enable precise TPR/FPR measurement impossible on larger benchmarks (SWE-bench: 29.6% label errors xia2025swebench). We trade size for quality. Testing 200+ samples would tighten intervals to ±7%.
@@ 7.5 Deployment Implications
Security-critical: Our 87.5% on security bugs with 100% compound detection works for finance, healthcare, infrastructure. Deploy as pre-commit gate blocking security HIGH (Algorithm 1). The 50% FPR is acceptable when security breach costs millions vs. developer time reviewing false alarms.
@@ 7.6 Future Work
More Compounds. We detect 4 pairs. Security literature has 100+ attack chains (MITRE ATT&CK, OWASP). Expand $E$ and test 3-way compounds (injection + credentials + crypto). $O(|V|^{2})$ extends to $O(|V|^{3})$, feasible for $|V|<20$.

### v3

# Multi-Agent Code Verification via Information Theory

@@ Abstract
LLMs generate buggy code: 29.6% of SWE-bench “solved” patches fail, 62% of BaxBench solutions have vulnerabilities, and existing tools only catch 65% of bugs with 35% false positives. We built CodeX-Verify, a multi-agent system that uses four specialized agents to detect different types of bugs. We prove mathematically that combining agents with different detection patterns finds more bugs than any single agent when the agents look for different problems, using submodularity of mutual information under conditional independence. Measuring agent correlation of $\rho=0.05$$0.25$ confirms they detect different bugs. Testing on 99 code samples with verified labels shows our system catches 76.1% of bugs, matching the best existing method (Meta Prompt Testing: 75%) while running faster and without test execution. We tested all 15 agent combinations and found that using multiple agents improves accuracy by 39.7 percentage points (from 32.8% to 72.4%) compared to single agents, with diminishing returns of +14.9pp, +13.5pp, and +11.2pp for agents 2, 3, and 4, validating our theoretical model. The best two-agent combination (Correctness + Performance) reaches 79.3% accuracy. Testing on 300 real patches from Claude Sonnet 4.5 runs in under 200ms per sample, making this practical for production use.
Keywords: Multi-agent systems, Code verification, LLM-generated code, Information theory
@@ 1 Introduction
Our Approach. We built CodeX-Verify, a system that runs four specialized agents in parallel: Correctness (logic errors, edge cases, exception handling), Security (OWASP Top 10, CWE patterns, secrets), Performance (algorithmic complexity, resource leaks), and Style (maintainability, documentation). Each agent looks for different bug types. We prove that combining agents finds more bugs than any single agent using submodularity of mutual information under conditional independence: $I(A_{1},A_{2},A_{3},A_{4};B)>\max_{i}I(A_{i};B)$. Measuring how often our agents agree shows correlation $\rho=0.05$$0.25$, confirming they catch different bugs.
@@
    Mathematical proof via submodularity of mutual information that combining agents with conditionally independent detection patterns finds more bugs than any single agent: $I(A_{1},\ldots,A_{n};B)>\max_{i}I(A_{i};B)$ when multiple agents are informative. Measured agent correlation $\rho=0.05$$0.25$ shows low redundancy.
@@
    Proof via submodularity that marginal information gains decrease monotonically (diminishing returns), validated by measured gains of +14.9pp, +13.5pp, +11.2pp for agents 2, 3, 4.
@@
    Comprehensive ablation testing all 15 agent combinations on a 29-sample hand-curated subset, showing multi-agent improves accuracy by 39.7 percentage points over single agents (32.8% → 72.4%), with best pair (Correctness + Performance) achieving 79.3%.
@@
    System achieving 76.1% TPR (matching state-of-the-art Meta Prompt Testing at 75%) with ¡200ms latency via static analysis. Dataset of 99 samples with verified labels released open-source.
@@ 2.1 LLM Code Generation and Verification
Wang and Zhu wang2024metamorphic propose Meta Prompt Testing: generate code variants with paraphrased prompts and detect bugs by checking if outputs differ. This achieves 75% TPR with 8.6% FPR on HumanEval. It requires test execution infrastructure and misses security vulnerabilities (SQL injection produces consistent outputs despite being exploitable) and quality issues that don’t affect outputs. AutoReview autoreview2025 uses three LLM agents (detector, locator, repairer) to find security bugs, improving F1 by 18.72% on ReposVul, but only checks security, not correctness or performance. We differ by: (1) proving mathematically why multi-agent works using submodularity of mutual information, (2) testing all 15 agent combinations to validate the architecture, and (3) achieving competitive TPR (76
@@ 2.3 Static Analysis and Vulnerability Detection
We extend static analysis by: (1) coordinating multiple agents that check different bug types, (2) proving mathematically why combining analyzers works via submodularity of mutual information, and (3) empirically validating the architecture through comprehensive ablation testing.
@@ 2.4 Ensemble Learning and Information Theory
Sheyner et al. sheyner2002attack model multi-step network exploits as attack graphs (directed graphs of vulnerability chains). Later work bayesian2018attack adds Bayesian risk and CVSS scoring cvss2024. Attack graphs focus on network vulnerabilities (host compromise, privilege escalation) rather than code-level bug detection, which our work addresses through multi-agent static analysis.
@@ 3 Theoretical Framework
We prove why multi-agent code verification beats single-agent approaches and derive sample size requirements using information theory and PAC learning. Section 6 tests all theoretical predictions.
@@ 3.2 Why Multi-Agent Works
Assume agents $A_{1},\ldots,A_{n}$ are pairwise conditionally independent given bug status $B$: $I(A_{i};A_{j}|B)=0$ for $i\neq j$. Then $f(S)=I(A_{S};B)$ is monotone submodular (Krause & Guestrin, 2011), implying:
$I(A_{1},\ldots,A_{n};B)>I(A_{i};B)\text{ for any single }i$ (2)
when multiple agents are informative about $B$.
@@
Under pairwise conditional independence given $B$, mutual information $f(S)=I(A_{S};B)$ is monotone submodular (Krause & Guestrin, 2011). Monotonicity states $f(S\cup\{i\})\geq f(S)$ for all $S,i$. Applying this recursively: $I(A_{1},\ldots,A_{n};B)=f(\{A_{1},\ldots,A_{n}\})\geq f(\{A_{i}\})=I(A_{i};B)$ for any single agent $i$. When multiple agents are informative ($I(A_{j};B)>0$ for some $j\neq i$), submodularity gives strict inequality. ∎
Our agents check different bug types: Correctness (logic errors, exception handling), Security (injection, secrets), Performance (complexity, resource leaks), Style (maintainability). These represent largely non-overlapping bug categories, supporting the conditional independence assumption. Measured unconditional correlation $\rho=0.05$$0.25$ shows low redundancy, though we cannot directly verify $I(A_{i};A_{j}|B)=0$ without conditional correlation measurements.
Measured Results. Single agents achieve 17.2% to 75.9% accuracy. Combined, 4 agents achieve 72.4% (average across configs). Progressive improvement (32.8% → 47.7% → 61.2% → 72.4% for 1, 2, 3, 4 agents) confirms the theorem.
3.3 Diminishing Returns
Theorem 2 (Diminishing Returns via Submodularity).
Assume agents $A_{1},\ldots,A_{n}$ are conditionally independent given $B$ (Theorem 1). Then $f(S)=I(A_{S};B)$ is monotone submodular in $S$, implying for any $S\subseteq T$:
$I(A_{i};B\mid A_{S})\geq I(A_{i};B\mid A_{T})$ (3)
When agents are ordered by decreasing performance, marginal information gains decrease: $\Delta I_{k}\geq\Delta I_{k+1}$ where $\Delta I_{k}=I(A_{k};B\mid A_{1},\ldots,A_{k-1})$.
Proof.
Under conditional independence given $B$, mutual information $f(S)=I(A_{S};B)$ is monotone submodular in $S$ (Krause & Guestrin, 2011, Theorem 2.2). Submodularity is equivalent to diminishing returns: adding element $i$ to smaller set $S$ yields more information gain than adding to larger set $T\supseteq S$, i.e., $f(S\cup\{i\})-f(S)\geq f(T\cup\{i\})-f(T)$. For ordered agents where $S_{k}=\{A_{1},\ldots,A_{k}\}$, this directly gives $\Delta I_{k}=f(S_{k})-f(S_{k-1})\geq f(S_{k+1})-f(S_{k})=\Delta I_{k+1}$. ∎
Corollary 1 (Optimal Agent Count). Optimal $n^{*}$ is where marginal gain equals marginal cost. Our measurements: +14.9pp, +13.5pp, +11.2pp for agents 2, 3, 4 (Section 6.2). Extrapolating predicts agent 5 would add ¡10pp, confirming $n^{*}=4$ is near-optimal.
@@ 3.4 Weighted Aggregation
Weight Selection Heuristic. We set agent weights based on three factors: individual accuracy $p_{i}$, redundancy with other agents $\bar{\rho}_{i}=\frac{1}{n-1}\sum_{j\neq i}\rho_{ij}$, and domain criticality $\gamma_{i}$. Combining these heuristically:
$w_{i}\propto p_{i}\cdot(1-\bar{\rho}_{i})\cdot\gamma_{i}$ (4)
Higher-accuracy agents receive higher weight, but weight decreases if the agent is redundant (high $\bar{\rho}_{i}$). The criticality term $\gamma_{i}$ captures asymmetric costs: security bugs block deployment more than style issues, justifying higher security weight despite lower solo accuracy.
@@ 4.1 Architecture
CodeX-Verify runs a pipeline: code input $\to$ parallel agent execution $\to$ weighted result aggregation $\to$ deployment decision (Figure 1).
@@
              (Weighted score combination)
@@
Design: (1) Agents check different bug types ($\rho=0.05$$0.25$ correlation). (2) Run in parallel via asyncio, ¡200ms latency. (3) Combine weighted scores and make deployment decision.
@@ 4.3 Aggregation
Agents run in parallel via asyncio (150ms max vs. 450ms sequential). Aggregation: collect issues from all agents, adjust severities based on code context, merge duplicate detections, compute weighted score $S_{\text{sys}}=\sum_{i}w_{i}S_{i}$, and apply decision thresholds (Algorithm 1).
@@ 4.4 Decision Logic
2:  if critical issues detected then
@@ 5.3 Ablation Design
We test all 15 combinations: single agents (4), pairs (6), triples (4), full system (1). Each configuration tested on the 29 hand-curated mirror samples (computational cost: 15 × 29 = 435 evaluations). Main evaluation (Section 6.1) uses all 99 samples (29 mirror + 70 Claude-generated). Hypothesis: Theorem 1 predicts multi-agent beats single when correlation is low, with diminishing returns (Theorem 2). Marginal contribution: $\Delta_{i}=\mathbb{E}[\text{Acc}(\text{with }A_{i})]-\mathbb{E}[\text{Acc}(\text{without }A_{i})]$.
@@ 6.1 Main Evaluation Results
All baseline numbers from literature (different datasets/methodologies). Direct comparison would require running all systems on identical samples.
@@ 7.3 Limitations
Sample Size. $n=99$ gives ±9.1% confidence intervals, wider than ideal. PAC bound (Theorem 3) suggests $n\geq 127$ for $\epsilon=0.15$, so we’re below optimal. But our perfect labels (100% verified) enable precise TPR/FPR measurement impossible on larger benchmarks (SWE-bench: 29.6% label errors xia2025swebench). We trade size for quality. Testing 200+ samples would tighten intervals to ±7%. Ablation study used 29-sample subset for computational efficiency; expanding to all 99 samples would increase statistical power.
@@ 7.4 Deployment Implications
Security-critical: Our 87.5% detection on security vulnerabilities works for finance, healthcare, infrastructure. Deploy as pre-commit gate blocking security HIGH issues (Algorithm 1). The 50% FPR is acceptable when security breach costs millions vs. developer time reviewing false alarms.
@@ 7.5 Future Work
Operating Point Analysis. Our evaluation reports a single operating point (76% TPR, 50% FPR). Characterizing the full TPR-FPR tradeoff curve via threshold sweeps, ROC/PR curve analysis, and per-category calibration would enable deployment-specific operating point selection and cost-sensitive decision making.